Solve for x and check the result.
(5/9)x + (-7) = (32/9)x + 8

Answer by stanbon(75887)

(5/9)x + (-7) = (32/9)x + 8
Multiply thru by 9 to get rid of the fractions:
5x -35 = 32x + 40
27x = -75
x = -75/27
x = -2.777777...
Cheers,
Stan H.